OpenStudy (anonymous):

You have four $1 bills, two $5 bills, five $10 bills, and five $20 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill. Find P($1, then $10).

OpenStudy (anonymous):

A. 9/39 B. 5/64 C. 3/80 D. 1/12

OpenStudy (anonymous):

is it B?

OpenStudy (anonymous):

i think it's b

OpenStudy (kropot72):

There are 16 bills of which 4 are $1. The probability of selecting a $1 on the first selection is 4/16. There would be 15 bills remaining after making the first choice. 5 of these 15 bills would be $10 if a $1 was already selected. Therefore the probability of selecting a $10 bill on the second selection is 5/15. The required probability is found by multiplying 4/16 by 5/15.\[P($1\ then\ $10)=\frac{4}{16}\times \frac{5}{15}=\frac{1}{4}\times \frac{1}{3}=you\ can\ calculate\]